# Improved subsample-and-aggregate via the private modified winsorized mean

### v1

@@ Appendix B Full description of subsample-and-aggregate parameter settings
  • For the estimator of Yu et al., [35]: $T$ was set to $\lfloor\log n\rfloor$ and other parameters were left at the default values, as per the code provided kindly from Yu et al., [35].
  • For $\tilde{\mu}^{\prime}$: $C=1$, $\eta=0$ and $\beta=1.001$.
  • For the estimator of Biswas et al., [4]: the upper bound on the covariance was set to $50\sqrt{d}=50\sqrt{6}$ and the starting center was set to the 0 vector. Also, $R=80\sqrt{6}$ and $\beta=0.1$. Three iterations were used in both the mean and covariance steps, with $\rho$ set to $1/3c,1/2c,1/c$ with $c=(1/2+1/3+1)/2$.
  • For the estimator of Liu et al., 2021a [25]: $\alpha=0.125$, $R=80\sqrt{6}$, and $\delta=0.01$ and other parameters were left at the default values, as per the code provided kindly from Liu et al., 2021a [25].
Appendix C Proof of Theorem 3.1
We now prove Theorem 3.1. Let $\tilde{X}=X-\mu$, $\xi_{q,X}$ be the (left continuous) $q$th quantile of the distribution of the random variable $X\sim F$, $\hat{\xi}_{q,X}$ be the (non-private) estimated $q$th quantile based on $X$, and $\tilde{\xi}_{q,X}$ the private estimate $q$th quantile based on a random sample of size $n$, whose observations have the same distribution as $X$.
Proof of Theorem 3.1.
@@
$\xi_{1-7\zeta/4}\leq\hat{\xi}_{1-3\zeta/2}\leq\xi_{1-5\zeta/4}\leq\xi_{1-3\zeta/4}\leq\hat{\xi}_{1-\zeta/2}\leq\xi_{1-\zeta/4}$ (C.1)
$\xi_{\zeta/4}\leq\hat{\xi}_{\zeta/2}\leq\xi_{3\zeta/4}\leq\xi_{5\zeta/4}\leq\hat{\xi}_{3\zeta/2}\leq\xi_{7\zeta/4}.$ (C.2)
@@
$\hat{\xi}_{\zeta/2}\leq\tilde{\xi}_{\zeta,Y^{\prime}}\leq\hat{\xi}_{3\zeta/2}\qquad\text{and}\qquad\hat{\xi}_{1-3\zeta/2}\leq\tilde{\xi}_{1-\zeta,Y^{\prime}}\leq\hat{\xi}_{1-\zeta/2},$ (C.3)
@@
The inequality (C.3) is implied by $\zeta/2\leq F_{n}(\tilde{\xi}_{\zeta,Y^{\prime}})\leq\frac{3}{2}\zeta$ and $21-\frac{3}{2}\zeta\leq F_{n}(\tilde{\xi}_{1-\zeta,Y^{\prime}})\leq 1-\zeta/$ 2, where one may recall that $F_{n}(x)=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\leq x\right\}/n$ for any $x\in\mathbb{R}$. Note that these inequalities are implied by the statements “$F_{n}(\tilde{\xi}_{\zeta})$ is within $\zeta/2$ of $\zeta$”, and that “$F_{n}(\tilde{\xi}_{1-\zeta})$ is within $\zeta/2$ of $1-\zeta$”, respectively. That is, taking $q=\zeta$ and $t=\zeta/2$, we are concerned with the quantities $|q-F_{n}(\tilde{\xi}_{q})|<t$ and $|1-q-F_{n}(\tilde{\xi}_{1-q})|<t$. To prove that these inequalities hold with high probability, we can apply Lemma E.1, once taking $q=\zeta$ and $t=\zeta/2$ and again with $q=1-\zeta$ and $t=\zeta/2$.
In order to apply Lemma E.1, we must have then that three inequalities hold. Namely, $\zeta<1$, and
$\beta$ $\leq\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1},\text{and}$ (C.4)
$\beta$ $\leq\frac{\hat{\xi}_{1-\zeta/2}-\ell+1}{\hat{\xi}_{1-3\zeta/2}-\ell+1}.$ (C.5)
First, $\zeta<1$ holds by Condition 2. For (C.4), note that (C.2) directly implies
$\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1}$ $=1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\hat{\xi}_{3\zeta/2}+1}>1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\xi_{5\zeta/4}+1}>1+\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}>1+b_{n}.$
Similarly, for (C.5), (C.1) directly implies that
@@
Condition 3 directly says that $\beta<1+b_{n}$, giving (C.4) and (C.5). Thus, we can apply Lemma E.1. Applying Lemma E.1 gives
$\Pr\left(\left|\zeta-F_{n}(\tilde{\xi}_{\zeta})\right|>\zeta/2\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2),\quad\text{and}$
$\Pr\left(\left|1-\frac{\zeta}{2}-F_{n}(\tilde{\xi}_{1-\zeta/2})\right|>\frac{\zeta}{2}\right)$ $\leq\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2).$
As a result, on $E$, (C.3) fails to hold with probability at most,
$\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)+\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
@@
The final inequality follows applying (C.1) and (C.2) and Condition 2. Consider that, since $\varepsilon_{1}\wedge\varepsilon_{2}\geq 3/56$ and
@@
$\leq 2\left(\frac{\log((u-\ell)+1)}{\log\beta}+1\right)\exp\left(-\frac{n\frac{112}{3n}\log\left(\frac{32}{\delta}\left[\frac{\beta(u-\ell)}{\beta-1}\vee 1\right]\right)(3/56)}{2}\right)$
$\leq 2\left(\frac{(u-\ell)\beta}{\beta-1}+1\right)\frac{\delta}{32}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}\leq\frac{\delta}{8}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}=\frac{\delta}{8}.$
Thus, combining the preceding statements yields that, on $E$, we have that (C.3) holds with probability at least
@@
The next step is to show that $E$ occurs, i.e., the inequalities (C.1) and (C.2) hold, with high probability. Call the event on which the inequalities (C.3) hold $A_{1}$. The inequalities (C.1) and (C.2) when combined with (C.3) imply,
@@
The next portion of the proof is similar to that of Theorem 1 of Lugosi and Mendelson, [28], however some steps are different, as that proof contains an error, see Appendix F. Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\})\text{ and }\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, [28]. To this end, observe that $\Pr\left(\tilde{X}\geq\xi_{1-7\zeta/4,\tilde{X}}\right)=7\zeta/4$. Recall that $\tilde{X}=X-\mu$. Define $W_{k-}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\leq\mu+\xi_{k,\tilde{X}}\right\}$ and $W_{k+}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\geq\mu+\xi_{k,\tilde{X}}\right\}$. Then, note that
@@
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(k-a)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(k-a)/3}\right)=\exp\left(-\frac{n\zeta(k-a)^{2}/2}{k(1-k\zeta)+(k-a)/3}\right).$
@@
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(a^{\prime}-k)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(a^{\prime}-k)/3}\right)=\exp\left(-\frac{n\zeta(a^{\prime}-k)^{2}/2}{k(1-k\zeta)+(a^{\prime}-k)/3}\right).$
@@
$W_{1}\geq\frac{17}{8}\zeta n\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-4\zeta,\tilde{X}}\right\}\geq\frac{17}{8}\zeta n-2\eta n\geq\frac{17}{8}\zeta n-n\zeta/8\geq 2n\zeta,$
$W_{2}\geq(1-3\zeta/8)n\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-\zeta/4,\tilde{X}}\right\}\geq(1-3\zeta/8)n-2\eta n\geq n(1-\zeta/2).$
@@
These inequalities state that, on $A_{2}$, the number of centered points in the corrupted sample above $\xi_{1-4\zeta,\tilde{X}}$ is at least $2n\zeta$. In addition, on $A_{2}$, observe that the number of centered points in the corrupted sample below $\xi_{1-\zeta/4,\tilde{X}}$ is at least $n(1-\zeta/2)$. As a result, (C.1) holds on $A_{2}$. An analogous argument yields (C.2) on $A_{2}$. Therefore, $E\subset A_{2}$.
@@
$\begin{split}1-\frac{\delta}{8}-\frac{\delta}{8}=1-\frac{\delta}{4}.\end{split}$ (C.6)
@@
$\leq\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})+\frac{1}{n}\sum_{i=1}^{n}[\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})].$
@@
$\hskip 150.00023pt+\mathbb{E}(\tilde{X}\mathbbm{1}\left\{\xi_{4\zeta,\tilde{X}}<\tilde{X}<\xi_{1-\zeta/4,\tilde{X}}\right\})$
@@
$\hskip 150.00023pt-\mathbb{E}(\tilde{X}(\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\}+\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\}))$
$=\mathbb{E}((\xi_{4\zeta,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\})+\mathbb{E}((\xi_{1-\zeta/4,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\})$
@@
$\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})=\mu-|\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})|+|\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})|$ $\geq\mu-\mathcal{E}(\zeta/4,X).$
@@
we have $t\leq t^{*}$ and thus $P(\overline{\omega}\geq t^{*})\leq P(\overline{\omega}\geq t)\leq p$. Taking $p=\delta/4$, combined with (C.6) and the assumption that $\delta\geq 4e^{-n}$, yields that, with probability at least $21-\delta/4-\delta/4=1-\delta/$ 2, there exists universal constants $c,c^{\prime}>0$ such that
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{\zeta/4,X},\xi_{1-\zeta/4,X}}(X_{i})$ $\leq\mu+\mathcal{E}(4\zeta,X)+c\left(\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}+\frac{2\mathcal{E}(\zeta/4,X)\log(4/\delta)}{3n}\right)$
$\leq\mu+c^{\prime}\mathcal{E}(\zeta/4,X)+c\left(\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}\right).$
@@
following the argument in Lugosi and Mendelson, [28]. Using Chebyshev’s inequality, note that the maximal gap for a single term in the summation,
@@
Combining this with (C.6), the preceding tail argument, and the fact that $\Pr\left(Z_{1}\geq 1\right)=e^{-n}\leq\delta/4$, we have that with probability at least $1-\delta,$
@@
$1/4n$ $=\Pr\left(X\geq\xi_{1-1/4n,X}\right)\leq\Pr\left(|X|\geq\xi_{1-1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1-1/4n,X}^{2}}\implies\xi_{1-1/4n,X}\lesssim\sigma\sqrt{n}.$
@@
$1/4n$ $=\Pr\left(X\leq\xi_{1/4n,X}\right)\leq\Pr\left(|X|\geq-\xi_{1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1/4n,X}^{2}}\implies-\xi_{1/4n,X}\lesssim\sigma\sqrt{n}.$

### v3

@@ Appendix B Full description of subsample-and-aggregate parameter settings
  • For the estimator of Yu et al., (2024): $T$ was set to $\lfloor\log n\rfloor$ and other parameters were left at the default values, as per the code provided kindly from Yu et al., (2024).
  • For $\tilde{\mu}_{p}^{\prime}$: $C=1$, $\eta=0$ and $\beta=1.001$.
  • For the estimator of Biswas et al., (2020): the upper bound on the covariance was set to $50\sqrt{d}=50\sqrt{6}$ and the starting center was set to the 0 vector. Also, $R=80\sqrt{6}$ and $\beta=0.1$. Three iterations were used in both the mean and covariance steps, with $\rho$ set to $1/3c,1/2c,1/c$ with $c=(1/2+1/3+1)/2$.
  • For the estimator of Liu et al., 2021a : $\alpha=0.125$, $R=80\sqrt{6}$, and $\delta=0.01$ and other parameters were left at the default values, as per the code provided kindly from Liu et al., 2021a .
Appendix C Proof of Theorem 1
We now prove Theorem 1. Let $\tilde{X}=X-\mu$, $\xi_{q,X}$ be the (left continuous) $q$th quantile of the distribution of the random variable $X\sim F$, $\hat{\xi}_{q,X}$ be the (non-private) estimated $q$th quantile based on $X$, and $\tilde{\xi}_{q,X}$ the private estimate $q$th quantile based on a random sample of size $n$, whose observations have the same distribution as $X$.
Proof of Theorem 1.
@@ Appendix C Proof of Theorem 1
$\xi_{1-7\zeta/4}\leq\hat{\xi}_{1-3\zeta/2}\leq\xi_{1-5\zeta/4}\leq\xi_{1-3\zeta/4}\leq\hat{\xi}_{1-\zeta/2}\leq\xi_{1-\zeta/4}$ (4)
$\xi_{\zeta/4}\leq\hat{\xi}_{\zeta/2}\leq\xi_{3\zeta/4}\leq\xi_{5\zeta/4}\leq\hat{\xi}_{3\zeta/2}\leq\xi_{7\zeta/4}.$ (5)
@@
$\hat{\xi}_{\zeta/2}\leq\tilde{\xi}_{\zeta,Y^{\prime}}\leq\hat{\xi}_{3\zeta/2}\qquad\text{and}\qquad\hat{\xi}_{1-3\zeta/2}\leq\tilde{\xi}_{1-\zeta,Y^{\prime}}\leq\hat{\xi}_{1-\zeta/2},$ (6)
@@
The inequality (6) is implied by $\zeta/2\leq F_{n}(\tilde{\xi}_{\zeta,Y^{\prime}})\leq\frac{3}{2}\zeta$ and $21-\frac{3}{2}\zeta\leq F_{n}(\tilde{\xi}_{1-\zeta,Y^{\prime}})\leq 1-\zeta/$ 2, where one may recall that $F_{n}(x)=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\leq x\right\}/n$ for any $x\in\mathbb{R}$. Note that these inequalities are implied by the statements “$F_{n}(\tilde{\xi}_{\zeta})$ is within $\zeta/2$ of $\zeta$”, and that “$F_{n}(\tilde{\xi}_{1-\zeta})$ is within $\zeta/2$ of $1-\zeta$”, respectively. That is, taking $q=\zeta$ and $t=\zeta/2$, we are concerned with the quantities $|q-F_{n}(\tilde{\xi}_{q})|<t$ and $|1-q-F_{n}(\tilde{\xi}_{1-q})|<t$. To prove that these inequalities hold with high probability, we can apply Lemma 2, once taking $q=\zeta$ and $t=\zeta/2$ and again with $q=1-\zeta$ and $t=\zeta/2$.
In order to apply Lemma 2, we must have then that three inequalities hold. Namely, $\zeta<1$, and
$\beta$ $\leq\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1},\text{and}$ (7)
$\beta$ $\leq\frac{\hat{\xi}_{1-\zeta/2}-\ell+1}{\hat{\xi}_{1-3\zeta/2}-\ell+1}.$ (8)
First, $\zeta<1$ holds by Condition 2. For (7), note that (5) directly implies
$\frac{u-\hat{\xi}_{\zeta/2}+1}{u-\hat{\xi}_{3\zeta/2}+1}$ $=1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\hat{\xi}_{3\zeta/2}+1}$
$>1+\frac{\hat{\xi}_{3\zeta/2}-\hat{\xi}_{\zeta/2}}{u-\xi_{5\zeta/4}+1}$
$>1+\frac{\xi_{5\zeta/4}-\xi_{3\zeta/4}}{u-\xi_{5\zeta/4}+1}>1+b_{n}.$
Similarly, for (8), (4) directly implies that
@@
Condition 3 directly says that $\beta<1+b_{n}$, giving (7) and (8). Thus, we can apply Lemma 2. Applying Lemma 2 gives
$\Pr\left(\left|\zeta-F_{n}(\tilde{\xi}_{\zeta})\right|>\zeta/2\right)$ $\leq\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2),$
and
$\Pr\left(\left|1-\frac{\zeta}{2}-F_{n}(\tilde{\xi}_{1-\zeta/2})\right|>\frac{\zeta}{2}\right)$ $\leq\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)$
×$\hskip 50.00008pt\times\exp\left(-\frac{n\zeta(\epsilon_{1}\wedge\epsilon_{2})}{2}\right).$
As a result, on $E$, (6) fails to hold with probability at most,
$\left(\frac{\log(u-\hat{\xi}_{3\zeta/2}+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
$\hskip 100.00015pt+\left(\frac{\log(\hat{\xi}_{1-3\zeta/2}-\ell+1)}{\log\beta}+1\right)\exp(-n\zeta(\epsilon_{1}\wedge\epsilon_{2})/2)$
@@
The final inequality follows applying (4) and (5) and Condition 2. Consider that, since $\varepsilon_{1}\wedge\varepsilon_{2}\geq 3/56$ and
@@
$\indent\leq 2\left(\frac{\log((u-\ell)+1)}{\log\beta}+1\right)\exp\left(-\frac{n\frac{112}{3n}\log\left(\frac{32}{\delta}\left[\frac{\beta(u-\ell)}{\beta-1}\vee 1\right]\right)(3/56)}{2}\right)$
$\indent\leq 2\left(\frac{(u-\ell)\beta}{\beta-1}+1\right)\frac{\delta}{32}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}$
$\indent\leq\frac{\delta}{8}\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)\left(\frac{\beta(u-\ell)}{\beta-1}\vee 1\right)^{-1}$
$\indent=\frac{\delta}{8}.$
Thus, combining the preceding statements yields that, on $E$, we have that (6) holds with probability at least
@@
The next step is to show that $E$ occurs, i.e., the inequalities (4) and (5) hold, with high probability. Call the event on which the inequalities (6) hold $A_{1}$. The inequalities (4) and (5) when combined with (6) imply,
@@
The next portion of the proof is similar to that of Theorem 1 of Lugosi and Mendelson, (2021), however some steps are different, as that proof contains an error, see Appendix G. Namely, one step uses the inequalities
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2\tilde{X}}}(X)\leq\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\geq\xi_{1-\zeta/2,X}\right\}),$
and
$\mathbb{E}\phi_{\xi_{2\zeta,\tilde{X}},\xi_{1-\zeta/2,X}}(X)\geq-\mathbb{E}(\tilde{X}\mathbbm{1}\left\{X\leq\xi_{2\zeta,X}\right\}),$
which do not hold, see page 400 and 401 of Lugosi and Mendelson, (2021). To this end, observe that $\Pr\left(\tilde{X}\geq\xi_{1-7\zeta/4,\tilde{X}}\right)=7\zeta/4$. Recall that $\tilde{X}=X-\mu$. Define $W_{k-}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\leq\mu+\xi_{k,\tilde{X}}\right\}$ and $W_{k+}=\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}\geq\mu+\xi_{k,\tilde{X}}\right\}$. Then, note that
@@
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(k-a)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(k-a)/3}\right)$
$=\exp\left(-\frac{n\zeta(k-a)^{2}/2}{k(1-k\zeta)+(k-a)/3}\right).$
@@
$\leq\exp\left(-\frac{n^{2}\zeta^{2}(a^{\prime}-k)^{2}/2}{nk\zeta(1-k\zeta)+n\zeta(a^{\prime}-k)/3}\right)$
$=\exp\left(-\frac{n\zeta(a^{\prime}-k)^{2}/2}{k(1-k\zeta)+(a^{\prime}-k)/3}\right).$
@@
$W_{1}\geq\frac{17}{8}\zeta n$
$\hskip 20.00003pt\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-4\zeta,\tilde{X}}\right\}\geq\frac{17}{8}\zeta n-2\eta n\geq\frac{17}{8}\zeta n-n\zeta/8\geq 2n\zeta,$
$W_{2}\geq(1-3\zeta/8)n$
$\hskip 20.00003pt\implies\sum_{i=1}^{n}\mathbbm{1}\left\{Y_{i}^{\prime}\geq\mu+\xi_{1-\zeta/4,\tilde{X}}\right\}\geq(1-3\zeta/8)n-2\eta n\geq n(1-\zeta/2).$
@@
These inequalities state that, on $A_{2}$, the number of centered points in the corrupted sample above $\xi_{1-4\zeta,\tilde{X}}$ is at least $2n\zeta$. In addition, on $A_{2}$, observe that the number of centered points in the corrupted sample below $\xi_{1-\zeta/4,\tilde{X}}$ is at least $n(1-\zeta/2)$. As a result, (4) holds on $A_{2}$. An analogous argument yields (5) on $A_{2}$. Therefore, $E\subset A_{2}$.
@@
$\begin{split}1-\frac{\delta}{8}-\frac{\delta}{8}=1-\frac{\delta}{4}.\end{split}$ (9)
@@
$\leq\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})$
$\hskip 20.00003pt+\frac{1}{n}\sum_{i=1}^{n}[\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{i})-\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})].$
@@
$\hskip 81.00014pt+\mathbb{E}(\tilde{X}\mathbbm{1}\left\{\xi_{4\zeta,\tilde{X}}<\tilde{X}<\xi_{1-\zeta/4,\tilde{X}}\right\})$
@@
$\hskip 81.00014pt-\mathbb{E}(\tilde{X}(\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\}+\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\}))$
$=\mathbb{E}((\xi_{4\zeta,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\leq\xi_{4\zeta,\tilde{X}}\right\})$
$\hskip 81.00014pt+\mathbb{E}((\xi_{1-\zeta/4,\tilde{X}}-\tilde{X})\mathbbm{1}\left\{\tilde{X}\geq\xi_{1-\zeta/4,\tilde{X}}\right\})$
@@
$\mathbb{E}\phi_{\xi_{4\zeta,X},\xi_{1-\zeta/4,X}}(X_{1})$ $=\mu-|\mathbb{E}(\xi_{1-\zeta/4,\tilde{X}}\wedge\tilde{X})|+|\mathbb{E}(\xi_{4\zeta,\tilde{X}}\vee\tilde{X})|$
$\geq\mu-\mathcal{E}(\zeta/4,X).$
@@
we have $t\leq t^{*}$ and thus $P(\overline{\omega}\geq t^{*})\leq P(\overline{\omega}\geq t)\leq p$. Taking $p=\delta/4$, combined with (9) and the assumption that $\delta\geq 4e^{-n}$, yields that, with probability at least $21-\delta/4-\delta/4=1-\delta/$ 2, there exists universal constants $c,c^{\prime}>0$ such that
$\frac{1}{n}\sum_{i=1}^{n}\phi_{\xi_{\zeta/4,X},\xi_{1-\zeta/4,X}}(X_{i})$ $\leq\mu+\mathcal{E}(4\zeta,X)$
$\hskip 20.00003pt+c\Bigg{(}\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}$
$\hskip 80.00012pt+\frac{2\mathcal{E}(\zeta/4,X)\log(4/\delta)}{3n}\Bigg{)}$
$\leq\mu+c^{\prime}\mathcal{E}(\zeta/4,X)$
$\hskip 40.00006pt+c\left(\sigma\sqrt{\frac{\log(4/\delta)}{n}}+\frac{2\xi_{1-\zeta/4,\tilde{X}}\log(4/\delta)}{3n}\right).$
@@
following the argument in Lugosi and Mendelson, (2021). Using Chebyshev’s inequality, note that the maximal gap for a single term in the summation,
@@
Combining this with (9), the preceding tail argument, and the fact that $\Pr\left(Z_{1}\geq 1\right)=e^{-n}\leq\delta/4$, we have that with probability at least $1-\delta,$
@@
$1/4n$ $=\Pr\left(X\geq\xi_{1-1/4n,X}\right)\leq\Pr\left(|X|\geq\xi_{1-1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1-1/4n,X}^{2}}$
$\hskip 60.00009pt\implies\xi_{1-1/4n,X}\lesssim\sigma\sqrt{n}.$
@@
$1/4n$ $=\Pr\left(X\leq\xi_{1/4n,X}\right)\leq\Pr\left(|X|\geq-\xi_{1/4n,X}\right)\leq\frac{\sigma^{2}}{\xi_{1/4n,X}^{2}}$
$\hskip 60.00009pt\implies-\xi_{1/4n,X}\lesssim\sigma\sqrt{n}.$